# IMAS2: Joint Agent Selection and Information-Theoretic Coordinated Perception In Dec-POMDPs

### v1

@@ 4.2. Policy Synthesis for Maximizing Information Gain
In Molloy and Nair (2023), the authors develop a method to compute an optimal policy that maximizes $I(X;Y,M_{\theta})$ for a single-agent POMDP. This method can be used for solving the inner maximization problem: First, we compute the single-agent POMDP for agent $j$ by fixing the selected agents’ policies $\bm{\pi}_{\mathcal{K}^{(i)}}$ computed from the previous iteration; second, the algorithm for trajectory estimation in single-agent POMDP Molloy and Nair (2023) can be applied. Note that the observations received by selected agents are included in the observation of the single-agent POMDP.
@@
The method in Shi et al. (2024) employs a policy gradient approach. For agent $i$, Let $\{\pi_{i,\theta}|\theta_{i}\in\Theta\}$ be a class of parameterized stochastic policies. We denote the joint policy parameter for all agents as $\bm{\theta}$ and the corresponding joint policy as $\bm{\pi}_{\bm{\theta}}$. Let $Y(\theta_{j})$ be agent $j$’s observation under agent $j$’s policy parameterized by $\theta_{i}$ and $\bm{Y}(\bm{\theta}_{\mathcal{K}^{(i)}})$ is the joint observation of agents in $\mathcal{K}^{(i)}$ given the joint policy of selected agents.
@@
Notably, when evaluating the conditional entropy, we use the joint observation $\bm{y}$ including both the observations of selected agents and the agent $j$’s observation under the current policy $\theta_{j}^{\tau}$. When we are evaluating the gradient of the log probability of observations, we only use the observations of agent $j$ because the other selected agents’ policies are fixed and independent from $\theta_{j}$. We have the following result from Shi et al. (2024):
@@
where $o_{j,0:t-1}$ is the local observation of agent $j$. The policy gradient from minimizing $H(Z|\bm{Y}(\bm{\theta}))$ is derived analogously (see Shi et al. (2024) for technical details.)
@@ 5. Experiment
Partial observable multi-agent system
The underlying environment state $s_{e}$ is partially observed by a multi-agent system, which is a directed sensor network. The sensor can change its limited field of view by choosing different angles of viewing. The sensing ranges of these sensors are illustrated in the top two panels of Figure 2. Each sensor has four possible sensing actions: NW, NE, SW, and SE. A sensor may detect the robot if the robot lies within the chosen sensing direction. We consider a weak sensor scenario. The detection process is stochastic: if the robot is within range, it is detected with probability $0.5$; otherwise, with probability $0.5$, the sensor receives a null observation ($n$). When the robot is outside the sensing range, the sensor also receives a null observation with probability $1$. Note that this multi-agent system, including the uncontrollable environment agent, can be modeled as a Dec-POMDP satisfying Assumption 2.
Inference objective
@@
For each selected sensor, we employ a long short-term memory (LSTM) neural network-based policy that directly processes observation sequences to action decisions. We implement Algorithm 1 in which the individual sensor’s optimal sensing strategy is computed using the policy gradient method in Shi et al. (2024). Figure 3(a) illustrates the convergence trend in a setting where the dynamics are deterministic and the sensor range is small  [^1] We sample $M=100$ trajectories and set the horizon $T=15$ for each iteration. The fixed learning rate of the policy gradient algorithm is set to be $0.001$. The hidden dimensions of all layers are set to be $64$. We run $N=1000$ iterations for each sensor on the 12th Gen Intel(R) Core(TM) i7-12700; the average time consumed for one iteration is $1.3$ seconds..
@@
IPG 0.525 75.5% 5.59 s
Table 1. Baseline comparison between IMAS2 and IPG.
Baseline Comparison. Existing approaches cannot directly address the joint problem of sensing-agent selection and decentralized active perception. Therefore, we compare our proposed IMAS2 algorithm against a variant of the Independent Policy Gradient (IPG) method Daskalakis et al. (2020). Since the IPG method lacks a mechanism for selecting optimal sensor subsets, we provide it with a fixed, randomly chosen group of sensors ($1,3,5,6,7$) from the nine available candidates. Both methods are evaluated under the stochastic environment with a large sensor range.
As shown in Table 1, the proposed IMAS2 algorithm achieves a lower conditional entropy ($0.493$) compared to IPG ($0.525$), indicating improved estimation accuracy. Actually, IMAS2 gives a higher inference accuracy ($86.0\%$) compared to IPG ($75.7\%$). Moreover, IMAS2 converges substantially faster, requiring only $1.5\,\text{s}$ per iteration—approximately $3.6$ times faster than the IPG baseline ($5.5\,\text{s}$). These results demonstrate that IMAS2 effectively balances computational efficiency and information gain in decentralized active perception tasks.
@@ 6. Conclusion
Appendix A Proofs of Lemmas

### v2

@@ 4.2. Policy Synthesis for Maximizing Information Gain
In Molloy2023smoother, the authors develop a method to compute an optimal policy that maximizes $I(X;Y,M_{\theta})$ for a single-agent POMDP. This method can be used for solving the inner maximization problem: First, we compute the single-agent POMDP for agent $j$ by fixing the selected agents’ policies $\bm{\pi}_{\mathcal{K}^{(i)}}$ computed from the previous iteration; second, the algorithm for trajectory estimation in single-agent POMDP Molloy2023smoother can be applied. Note that the observations received by selected agents are included in the observation of the single-agent POMDP.
@@
The method in shi2024active employs a policy gradient approach. For agent $i$, Let $\{\pi_{i,\theta}|\theta_{i}\in\Theta\}$ be a class of parameterized stochastic policies. We denote the joint policy parameter for all agents as $\bm{\theta}$ and the corresponding joint policy as $\bm{\pi}_{\bm{\theta}}$. Let $Y(\theta_{j})$ be agent $j$’s observation under agent $j$’s policy parameterized by $\theta_{i}$ and $\bm{Y}(\bm{\theta}_{\mathcal{K}^{(i)}})$ is the joint observation of agents in $\mathcal{K}^{(i)}$ given the joint policy of selected agents.
@@
Notably, when evaluating the conditional entropy, we use the joint observation $\bm{y}$ including both the observations of selected agents and the agent $j$’s observation under the current policy $\theta_{j}^{\tau}$. When we are evaluating the gradient of the log probability of observations, we only use the observations of agent $j$ because the other selected agents’ policies are fixed and independent from $\theta_{j}$. We have the following result from shi2024active:
@@
where $o_{j,0:t-1}$ is the local observation of agent $j$. The policy gradient from minimizing $H(Z|\bm{Y}(\bm{\theta}))$ is derived analogously (see shi2024active for technical details.)
@@ 5. Experiment
[…]
Comparison of convergence, entropy, and inference performance under different sensing conditions.
Partial observable multi-agent system
The underlying environment state $s_{e}$ is partially observed by a multi-agent system, which is a directed sensor network. The sensor can change its limited field of view by choosing different angles of viewing. The sensing ranges of these sensors are illustrated in the top two panels of Figure 2. Each sensor has four possible sensing actions: NW, NE, SW, and SE. A sensor may detect the robot if the robot lies within the chosen sensing direction. If the robot is detected, the sensor will receive the exact robot location. We consider a weak sensor scenario. The detection process is stochastic: if the robot is within range, it is detected with probability $0.5$; otherwise, with probability $0.5$, the sensor receives a null observation ($n$). When the robot is outside the sensing range, the sensor also receives a null observation with probability $1$. Note that this multi-agent system, including the uncontrollable environment agent, can be modeled as a Dec-POMDP satisfying Assumption 2.
Inference objective
@@
For each selected sensor, we employ a long short-term memory (LSTM) neural network-based policy that directly processes observation sequences to action decisions. We implement Algorithm 1 in which the individual sensor’s optimal sensing strategy is computed using the policy gradient method in shi2024active. Figure 3(a) illustrates the convergence trend in a setting where the dynamics are deterministic and the sensor range is small  [^2] We sample $M=100$ trajectories and set the horizon $T=15$ for each iteration. The fixed learning rate of the policy gradient algorithm is set to be $0.001$. The hidden dimensions of all layers are set to be $64$. We run $N=1000$ iterations for each sensor on the 12th Gen Intel(R) Core(TM) i7-12700; the average time consumed for one iteration is $1.3$ seconds. The code is available at https://github.com/AronYoung414/multi-agent-active-perception-grid-world..
@@
Fixed Selector (IPG) 0.525 75.5% 7.62 s
Random Selector (IPG) 0.558 70.7% 7.63 s
Visibility-Based Selector (IPG) 0.502 84.1% 7.63 s
Baseline Comparison. Existing approaches cannot directly address the joint problem of sensing-agent selection and decentralized active perception. The sensing approaches in albrecht2017reasoning; shafipour2022online employ Bayesian inference mechanisms similar to those used in IMAS2. However, these works focus on passive observation models and do not design or optimize active sensing policies that control how information is gathered. And the key challenge is that our objective is mutual information, rather than a cumulative reward/value function. Consequently, standard value-function-based MARL algorithms (e.g., MADDPG lowe2017multi, MAPPO yu2021mappo) are not directly applicable, and classical Dec-POMDP solvers require an explicit reward structure or a belief-based value function. We therefore compare against a variant of the Independent Policy Gradient (IPG) method Daskalakis2020ipg, which is a decentralized gradient-based method that optimizes each agent’s policy directly via policy gradients without relying on a centralized value function. Since IPG does not include a mechanism for selecting an optimal sensor subset, we provide it with a fixed chosen group of sensors ($1,3,5,6,7$), randomly chosen group ($8,5,4,6,1$) of sensors, and visibility-based selector (The set ($0,2,4,5,8$) covers the most area) from the nine available candidates. Both methods are evaluated under the stochastic environment with a large sensor range.
As shown in Table 1, the proposed IMAS2 algorithm achieves a lower conditional entropy ($0.493$) compared to IPG ($0.525$, $0.558$, $0.502$), indicating improved estimation accuracy. Actually, IMAS2 gives a higher inference accuracy ($86.0\%$) in the test environment compared to IPG ($75.5\%$, $70.7\%$, $84.1\%$). Moreover, IMAS2 converges substantially faster, requiring only $1.5\,\text{s}$ per iteration—approximately $5.06$ times faster than the IPG baseline ($7.6\,\text{s}$). These results demonstrate that IMAS2 effectively balances computational efficiency and information gain in decentralized active perception tasks.
@@ 6. Conclusion
{acks}
Research was sponsored by the Army Research Laboratory under Cooperative Agreement Number W911NF-25-2-0045 and by the Army Research Office under Award Number W911NF-22-1-0166. ARO, as the Federal awarding agency, reserves a royalty-free, nonexclusive and irrevocable right to reproduce, publish, or otherwise use this software for Federal purposes, and to authorize others to do so in accordance with 2 CFR 200.315(b). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government.
Appendix A Proofs of Propositions, Lemmas, and Theorems
@@
Proof of Lemma 4 Let $\Omega\coloneqq\{Y_{i},i\in\mathcal{N}\}$ be the set of observations for all agents. $Y_{A}\subset Y_{B}\subset\Omega$, and $Y_{j}\in\Omega\setminus Y_{B}$.
$I(X;Y_{B})-I(X;Y_{A})=I(X;Y_{B}\setminus Y_{A}|Y_{A})$
Since conditional mutual information is always non-negative Cover2006EIT, $I(X;Y_{B})-I(X;Y_{A})\geq 0$ and thus the function $g(\cdot)$ is monotone. To show $g(\cdot)$ is submodular, we need to show that
$I(X;Y_{A}\cup\{Y_{j}\})-I(X;Y_{A})\geq I(X;Y_{B}\cup\{Y_{j}\})-I(X;Y_{B}).$
And based on the chain rule of mutual information,
$I(X;Y_{A}\cup\{Y_{j}\})-I(X;Y_{A})=I(X;Y_{j}|Y_{A}).$
$H(Y_{j}|Y_{A})\geq H(Y_{j}|Y_{B})$
$H(Y_{j}|Y_{A})-H(Y_{j}|X)\geq H(Y_{j}|Y_{B})-H(Y_{j}|X)$
$H(Y_{j}|Y_{A})-H(Y_{j}|X,Y_{A})\geq H(Y_{j}|Y_{B})-H(Y_{j}|X,Y_{B})$
$I(Y_{j};X|Y_{A})\geq I(Y_{j};X|Y_{B})$
$I(X;Y_{j}|Y_{A})\geq I(X;Y_{j}|Y_{B}),$
where the third step is due to Lemma 3 and the last step is because conditional mutual information is symmetric.
@@
Proof of Lemma 9 From (6), the mutual information between the secret $Z$ and the observations $Y_{A}$ can be written as
$I(Z;Y_{A})=I(X_{e};Y_{A})-H(X_{e})+H(Z)+H(X_{e}\mid Z,Y_{A}).$ (11)
Rearranging (11), we obtain
$I(X_{e};Y_{A})-I(Z;Y_{A})=H(X_{e})-H(Z)-H(X_{e}\mid Z,Y_{A}).$
Since $Z$ is a deterministic function of $X_{e}$, we have $H(Z\mid X_{e})=0$. By expanding the conditional mutual information,
$I(X_{e};Y_{A}\mid Z)=H(X_{e}\mid Z)-H(X_{e}\mid Z,Y_{A})=H(X_{e})-H(Z)-H(X_{e}\mid Z,Y_{A}).$
Therefore,
$I(Z;Y_{A})=I(X_{e};Y_{A})-I(X_{e};Y_{A}\mid Z).$
Define
$\epsilon\coloneqq\max_{A}\frac{I(X_{e};Y_{A}\mid Z)}{I(X_{e};Y_{A})}.$
Then for all $A$,
$I(Z;Y_{A})\geq(1-\epsilon)I(X_{e};Y_{A}).$
Moreover, by the data processing inequality and the fact that $Z$ is a function of $X_{e}$,
$I(Z;Y_{A})\leq I(X_{e};Y_{A}).$
Combining the above inequalities yields
$(1-\epsilon)I(X_{e};Y_{A})\leq I(Z;Y_{A})\leq I(X_{e};Y_{A}),$
which can be loosened to
$(1-\epsilon)I(X_{e};Y_{A})\leq I(Z;Y_{A})\leq(1+\epsilon)I(X_{e};Y_{A}).$
Since $I(X_{e};Y_{A})$ is monotone and submodular in $A$, this shows that $I(Z;Y_{A})$ is $\epsilon$-approximately submodular.
Proof of Proposition 10 Recall $(j^{\ast},\pi_{j^{\ast}})$ is the choice of agent and its policy selected at the $i$-th iteration of Algorithm 1. According to Algorithm 1,
$-f({\mathcal{K}}^{(i-2)},\bm{\pi}^{(i-2)})\big)$
$=$ $f({\mathcal{K}}^{(i-1)},\bm{\pi}^{(i-1)})-f({\mathcal{K}}^{(i-2)},\bm{\pi}^{(i-2)})=\Delta_{i}.$
Here, $(i)$ follows from the submodularity of the objective function: the marginal contribution of agent $j^{\ast}$ using policy $\pi_{j^{\ast}}$ with respect to $(\mathcal{K}^{(i-2)},\bm{\pi}^{(i-2)})$ is greater than or equal to its contribution with respect to $(\mathcal{K}^{(i-1)},\bm{\pi}^{(i-1)})$ under the same policy. And $(ii)$ is due to the greedy choice made at step i1i-1.
Proof of Theorem 11 For each $i\geq 1$, let the gap between the value under the optimal solution and the value under the solution returned by the $i$-th selection be $g_{i}\coloneqq f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})$.
$\overset{(ii)}{=}$ $k\Delta_{i+1}-\Delta_{i+2}$
$=$ $\;k\Delta_{i+2}-(k+1)\Delta_{i+2}+k\Delta_{i+1}$
$\overset{(iii)}{\leq}$ $k\Delta_{i+2}.$
Using successive gaps, $\Delta_{i+1}\geq\frac{1}{k}g_{i}$ and $g_{i+1}=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i+1)},\bm{\pi}^{(i+1)})=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i+1)},\bm{\pi}^{(i+1)})+f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})-f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})=g_{i}-\Delta_{i+1}\leq g_{i}-\frac{1}{k}g_{i}=(1-\frac{1}{k})g_{i}$. (This part of the proof is similar to the original proof in nemhauser1978analysis).
Consequentially, $g_{k}\leq(1-\frac{1}{k})^{k}g_{0}$ where $g_{0}=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})$. Using the exponential bound, $(1-\frac{1}{k})^{k}\leq e^{-1}$ for any $k\geq 1$. Equation (7) is then derived.